OpenStudy (anonymous):

For the given quadratic equation convert into vertex form, find the vertex, and find the value for x=6. Show your work. y=-2x^2 + 2x+ 2

OpenStudy (irishboy123):

-(1/2)y = x^2 - x - 1 (-1/2)y + 1 = x^2 - x (-1/2)y + 1 = (x - 1/2)^2 - 1/4 (-1/2)y = (x-1/2)^2 - 5/4 y = -2(x-1/2)^2 + 5/2

OpenStudy (anonymous):

Is that the vertex form?

OpenStudy (mathmate):

Yes, it is in vertex form.

OpenStudy (irishboy123):

important this is not so much the answer but that you dig what we are doing and how we went about it. we have rebased the equation around its vertex. that's what's important about this.

OpenStudy (anonymous):

Is this the right answer?